1.-(50) Find: Ro RG 15V Yes, Im and the region of geration k=60 mA/² = 20 IMR 12 / fa RO 200-12
Answer #1


: sol 1 oó ISV Ro Vax Iazo y S TIT :-) So the Given circuit is in saturation RD mi VID G Toiva = No Nosa Vas-VT region --> In =) ID = kn (vg-VJV ID = krivo - Vtv NG=VD 2 (2 from © 15-vb. kalvo-ut) 200 Is-Vo = (con) (200) (VD +V2Y=2V0V) 1SAVO = 12( 10